OpenStudy (anonymous):

find the 4 consecutive integers for -22 i did the work just making sure its right the work will be in the comments

OpenStudy (anonymous):

-7+-6+-5+-4=-22

OpenStudy (anonymous):

n+(n+1)+(n+2)+(n+3)=-22 4n+6=-22 -6 -6 4n=-28 4n/4=-28/4 n=-7 soooo... -7,-8,-9,-10

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

No because if you add those together you don't get -22 since 22 is negative you would do n-1 n-2 n-3 n-4 to get your outcomes

OpenStudy (anonymous):

oh so 7,8,9,-10

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so -7,-6,-5,-4

OpenStudy (anonymous):

Yes exactly. Your equation was correct andy you got n=-7. Then n+1=-6, n+2=-5 and n+3=-4 :)

OpenStudy (anonymous):

yay thank you